OpenStudy (anonymous):

true or false for angle Θ between 0 and π/2. if tanΘ=(1/3), then sinΘ = (√10)/10.

OpenStudy (anonymous):



OpenStudy (anonymous):

find the hypotenuse, then see if the sine gives you \[\frac{\sqrt{10}}{10}\]

OpenStudy (anonymous):

here is a hint: yes it does, because the hypotenuse is \[\sqrt{3^2+1^2}=\sqrt{10}\] and \[\frac{1}{\sqrt{10}}=\frac{\sqrt{10}}{10}\]

OpenStudy (aravindg):

myinmia u there???

OpenStudy (aravindg):

heloooooooooooo